# Entanglement engineering of optomechanical systems by reinforcement learning

### v1

@@ VI Discussion
Exploiting machine learning for controlling quantum information systems is becoming a promising field of research and is attracting increasing attention. We have developed a model-free deep-RL method for entanglement engineering. We demonstrated its superiority over benchmark quantum control methods in quantum optomechanical systems under WCM. The deep-RL agent sequentially interacts with one or multiple parallel quantum optomechanical environments, collects trajectories, and updates its policy to maximize the accumulated reward to create and stabilize the entanglement. Both linear and nonlinear interacting regimes between the photons in the optical cavity and the phonons associated with the mechanical oscillator in the cavity have been studied. In particular, for linear interactions, the PPO agent directly observes the WCM and delivers better performance compared with the benchmark Bayesian and random control methods in the framework of measurement-based feedback control. The performance of deep RL control is tolerant to randomness when initially the system is in some mixed state. For nonlinear interactions, both PPO and recurrent PPO agents have been tested, where the first was utilized to generate the time series of the target of the expected photon number, and the second one was employed to control entanglement according to an objective. Because of the high degree of randomness in the SME originating from ten stochastic collapse operators, only the observation of the expected photon number is feasible in the nonlinear interaction regime.
@@
More specifically, linear interactions can naturally limit the excitation in the energy levels, providing a mechanism to create the entangled Bell directly states under the premise of strong laser approximation in the red-detuned regime. A disadvantage is that its performance is sensitive to the coupling to the vacuum or thermal bath, even when the decay rate is small (e.g., $\kappa=0.01\,\omega_{m}$ with $G\in[-5,5]\omega_{m}$). This phenomenon is in fact quite common in quantum systems. For instance, in systems with magnon-photon coupling [90], steady Bell states can ideally be generated in the PT-broken phase without dissipation while the entanglement is reduced when the decay rate is non-negligible. Another issue with linear interactions is that the time scale associated with generating entangled Bell states [93] tends to be much shorter than the inverse of the coupling strength about the higher-order exceptional points in a system of coupled non-Hermitian qubits with energy loss while the maximum entanglement can only last for a short instant.
@@
In contrast, nonlinear interactions can create and stabilize entanglement and are more robust to the disturbance from the thermal bath even with a relatively large decay rate, e.g., $\kappa=0.1\,\omega_{m}$ with the strong coupling $g_{0}=0.2\,\omega_{m}$, so $g_{0}/\kappa=2>1$ to stipulate the nonlinear effect [66]. Potentially, systems with nonlinear coupling thus can outperform those with linear interactions. A caveat is that, in nonlinear optomechanical systems, there is limited experimentally accessible observation. In fact, the relationship between experimental observables other than the photon number and entanglement in open quantum optomechanical systems have not been well understood, rendering it challenging to choose a feasible observable to control entanglement. We have partially relied on the numerical model to create and stabilize entanglement, based on the numerical relation between entanglement and the expected photon number. Another difficulty is that the nonlinear interaction can readily excite the system to high quantum states, which we have overcome by designing a proper reward function in deep-RL.
@@
It is possible to exploit multi-agent reinforcement learning (MARL) through parallel computation to stabilize entanglement, where the agents exploit the decentralized structure of the task and share the information via communication. Saliently, if several agents fail in a multiagent system, the remaining agents can take over some of their tasks. In principle, our control framework can be extended to MARL for multi-mode entanglement engineering of a quantum black box.
Limitations and challenges of entanglement control.
For general physical system, a limitation lies in experimental observation design. Existing works used full fidelity, i.e., the overlap between the current and target quantum states, as their observation metric, ranging from the generation of two [94] and multi-qubit entangled states [95, 96] to specific many-body states [97, 98, 99] and single-particle quantum state engineering via deep RL [42, 43, 44, 45]. However, full fidelity observation is not universally applicable to experiments. Another limitation concerns quantum measurement effects. Projective measurement, where the quantum state completely collapses after observation, is commonly used, such as in the variational quantum circuit-based preparation of many-body ground states [100]. However, WCMs have not been extensively studied for entanglement control. Moreover, obtaining the relationship between entanglement and experimental observables is difficult. So far there have been no systematical methods to extract quantitative entanglement from experimental observation for arbitrary systems [101, 102, 103], in spite of some initial exploration for specific systems. For example, an entanglement criterion for non-Gaussian states in coupled harmonic oscillators was developed [101]. Under the strong laser approximation, a Bell inequality was tested with photon counting [102], and stationary entanglement for Gaussian states was inferred from the continuous measurement of light only [103].
For quantum control of optomechanical systems, most existing theoretical studies focused on Gaussian states or linear interaction regime [104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 109, 115, 116]. In addition, most existing control methods require the system model such as the pulse method [104, 105, 106, 107, 108], the time-continuous laser driven approach [109, 110], periodic modulations [111, 112, 113], optimal pulse protocol [114], linear quadratic-Gaussian (LQG) [109], and coherent feedback methods with auxiliary optical components [115, 116]. Moreover, there were works focusing on entanglement enhancement [114, 109, 115, 116].
The aim of our work is developing a deep-RL method to maintain non-Gaussian entanglement in optomechanical systems with linear or nonlinear interactions, using only photon number counting from WCM. To our knowledge, prior to our work, model-free feedback control to create and stabilize the entanglement of non-Gaussian states in nonlinear stochastic optomechanical systems had not been available. Our work overcomes some of the aforementioned limitations and challenges. In particular, concerning the experimental observation, in a related work [45], a PPO agent was used to create different Fock states and the superposition of a single cavity mode based on observing the density matrix and a fidelity based reward function. In contrast, the observable in our work is the photon current that is more experimentally accessible [69]. With respect to the issue of quantum measurement, we use WCM in real-time feedback control, taking into consideration the resulting quantum stochastic process [51, 62, 43]. Concerning the relationship between entanglement and photocurrent, in our model-free RL control, we identify the numerical relationship between entanglement and photocurrent.
We have demonstrated non-Gaussian state control with both linear and nonlinear interaction. The importance of studying non-Gaussian state control lies in the fact that, under the nonlinear quantum master equation resulting from WCM, the time-evolved quantum states are generally intrinsically non-Gaussian. Our deep RL control scheme is model-free [117], where policies or value functions are directly learned from the interactions with the quantum environment without any explicit model of this environment. This should be contrasted to the model-based deep RL methods [118], where a pre-built model of the environment for policy decision-making is needed. In our work, the RL agent utilizes only the observables it receives at each time step to inform its decisions, without relying on any explicit model of the underlying quantum system. In fact, the RL agent learns the optimal policy through direct interaction with the environment.
Experimental feasibility of RL based control.
Our control strategy was articulated by considering the current experimental capability. In particular, previous works studied the microwave regime of the spectrum in optomechanical systems [84, 83, 82] suggests the possibility of experimental implementation of our RL control scheme. It was demonstrated that the photon-number statistics of a microwave cavity mode can be detected using multiplexed photon number measurement [69]. In this method, the reduction in the reflection amplitude of the frequency comb is proportional to the photon-number distribution of the microwave cavity mode over the Fock bases, as detected by weak measurement [45]. The beam-splitter Hamiltonian [119, 120, 121] can be fabricated with a strong laser driven in the red-detuned regime and the nonlinear interaction in the experimental quantum optomechanical systems has been realized [122, 66, 123]. The control action required for our RL based scheme can be the amplitude or frequency of the input laser. Taken together, our work represents an initial effort of entanglement engineering in a quantum black box through model-free deep RL and measurement-based feedback control under the SME and WCM based on the current experimental feasibility.

### v2

@@ Discussion
Exploiting machine learning for controlling quantum information systems is becoming a promising research realm and is attracting increasing attention. We have developed a model-free deep-RL method for entanglement engineering. We demonstrated its superiority over benchmark quantum control methods in quantum optomechanical systems under WCM. The model-free deep-RL agent sequentially interacts with one or multiple parallel quantum optomechanical environments, collects trajectories, and updates its policy to maximize the accumulated reward to create and stabilize the entanglement. Both linear and nonlinear interacting regimes between the photons in the optical cavity and the phonons associated with the mechanical oscillator in the cavity have been studied. In particular, for linear interactions, the PPO agent directly observes the WCM photocurrent and delivers better performance compared with the benchmark Bayesian and random control methods in the framework of measurement-based feedback control. The performance of deep RL control is tolerant to randomness when initially the system is in some mixed state. For nonlinear interactions, both the model-free PPO and recurrent PPO agents have been tested, where the first was utilized to generate the time series of the target of the expected photon number, and the second one was employed to control entanglement according to an objective. Because of the high degree of randomness in the SME originating from ten stochastic collapse operators, only the observation of the expected photon number is feasible in the nonlinear interaction regime.
@@
More specifically, linear interactions can naturally limit the excitation in the energy levels, providing a mechanism to directly create the entangled Bell states under the premise of strong laser approximation in the red-detuned regime. A disadvantage is that its performance is sensitive to the coupling of the vacuum or thermal bath, even when the decay rate is small (e.g., $\kappa=0.01\,\omega_{m}$ with $G\in[-5,5]\,\omega_{m}$). This phenomenon is in fact quite common in quantum systems. For instance, in systems with magnon-photon coupling [87], steady Bell states can ideally be generated in the PT-broken phase without dissipation while the entanglement is reduced when the decay rate is not negligible. Another issue with linear interactions is that the time scale associated with generating entangled Bell states [89] tends to be much shorter than the inverse of the coupling strength about the higher-order exceptional points in a system of coupled non-Hermitian qubits with energy loss while the maximum entanglement can only last for a short instant.
@@
In contrast, nonlinear interactions can create and stabilize entanglement and are more robust to the disturbance from the vacuum bath even with a relatively large decay rate, e.g., $\kappa=0.1\,\omega_{m}$ with the strong coupling $g_{0}=0.2\,\omega_{m}$, so $g_{0}/\kappa=2>1$ to stipulate the nonlinear effect [61]. Potentially, systems with nonlinear coupling thus can outperform those with linear interactions. A caveat is that, in nonlinear optomechanical systems, there is limited experimentally accessible observation. In fact, the relationship between experimental observables and entanglement in nonlinear quantum optomechanical systems has not been well understood, rendering challenging to choose a feasible observable to control entanglement. We have partially relied on the numerical method to create and stabilize entanglement, based on the numerical relation between entanglement and the expected photon number discovered by the deep RL. Another difficulty is that the nonlinear interaction can readily excite the system to high quantum states, which we have overcome by designing a proper reward function.
A previous work [90] studied the acceleration of entanglement generation through feedback weak measurement for two qubits in a four-dimensional Hilbert space, where coupling to a vacuum or thermal bath was not taken into account, nor the interactions between the two qubits, and the control protocol required prior knowledge about the system such as the decoherence-free subspace. In addition, complete observation was needed to design the local Hamiltonian feedback to speed up entanglement. This is in fact a model-based approach. In another study [91], steady-state entanglement between two qubits was achieved using a continuous feedback control method, where the feedback protocol design was informed by a detailed model of the system’s dynamics. In contrast, our work creates and stabilizes a two-mode entangled state about a predetermined level of entanglement for both linear and nonlinear interactions via model-free reinforcement learning, with the respective dimensions of the Hilbert space being four and one hundred.
@@
Our work suggests the possibility of exploiting multi-agent RL through parallel computation to stabilize entanglement. The agents leverage the decentralized structure of the task and share information via communication. Saliently, if several agents fail in a multiagent system, the remaining agents can take over some of their tasks. In principle, our control framework can be extended to multi-agent RL for multi-mode entanglement engineering of a quantum black box.
Methods
Stochastic Master Equation
An experimental optomechanical system is effectively an open quantum system interacting with the vacuum bath under WCM with the operators [49, 29] $\hat{C}_{n}\equiv\sqrt{\eta}\hat{P}_{n}$, where $\hat{P}_{n}=|n\rangle\langle n|$ with $n=0,1$ (linear) or $n=0,1,\ldots,9$ (nonlinear) is the measurement operator on the Fock state and $\eta$ denotes the measurement rate. The quantum dynamics of this system are described by the stochastic master equation (SME) [92, 29, 93, 94, 51]:
$d\rho$ $=\frac{1}{i\hbar}[\tilde{H},\rho]dt+\mathcal{L}_{env}\;\rho dt$
$+\sum_{n}\mathcal{D}(\hat{C}_{n})\rho dt+\sum_{n}\mathcal{H}(\hat{C}_{n})\rho dW_{n},$ (2)
where the Hamiltonian is $\tilde{H}=\tilde{H}_{bs}$ or $\tilde{H}_{nl}$ and $\rho$ is a density operator in the Hilbert space. Under the Born-Markov approximation [95, 96], which requires the system-bath coupling to be weak and the correlation time of the bath to be much shorter than a characteristic timescale of system-bath interactions, the Markovian master equation, i.e., the first two terms in the right-hand side of Eq. (Stochastic Master Equation), has the Lindblad form [95]. At absolute zero temperature, the following environmental operator $\mathcal{L}_{env}\;\rho$ can be introduced to describe the coupling between the system and vacuum bath: $\mathcal{L}_{env}\;\rho=\kappa\mathcal{D}(\hat{a})\rho+\gamma\mathcal{D}(\hat{b})\rho$, where the cavity and oscillator modes are coupled to the vacuum bath with the strength $\kappa$ and $\gamma$, respectively [56]. The deep RL results in the Lindblad master equation with the nonlinear interaction are presented in Supplementary Note 5.
The WCM process described by the last two terms in the right-hand side of Eq. (Stochastic Master Equation) is nonlinear and Markovian in the unconditional master equation [51] in $\rho$. Under WCM, a Wiener process $dW$ with a Gaussian distribution [51] arises from the Gaussian-weighted projection over the eigenstates that allows the quantum information to be extracted continuously in the time domain, subject to stochastic disturbances in the last term of Eq. (Stochastic Master Equation) and quantum decoherence in the penultimate term of Eq. (Stochastic Master Equation). (Supplementary Note 3 in SI provides a detailed derivation of the SME.) The Lindblad operator $\mathcal{D}$ and the measurement superoperator $\mathcal{H}$ in Eq. (Stochastic Master Equation) are given by
$\mathcal{D}(\hat{A})\rho$ $\equiv\hat{A}\rho\hat{A}^{\dagger}-\frac{1}{2}(\hat{A}^{\dagger}\hat{A}\rho+\rho\hat{A}^{\dagger}\hat{A}),$
$\mathcal{H}(\hat{A})\rho$ $\equiv\hat{A}\rho+\rho\hat{A}^{\dagger}-\langle\hat{A}+\hat{A}^{\dagger}\rangle\rho,$
with $\langle\hat{A}\rangle\equiv\textnormal{Tr}[\hat{A}\rho]$. The two operators serve to weakly drive the quantum state into the corresponding eigenstates to some degree.
Implementation details of deep RL
For simulating the linear or nonlinear quantum optomechanical system described by Eq. (Stochastic Master Equation), we use the “taylor1.5” solver from the SME solver in the QuTip’s package [97] with the tolerance $\textnormal{tol}=10^{-6}$ and time step size $dt=0.01\,\omega_{m}^{-1}$. The measurement current is simulated with the “homodyne” method, and the custom environment is constructed by the open-source platform OpenAI-Gym [98]. For RL simulations, we construct the PPO agent [48] by “stable-baselines3” [99] in the A2C [100] settings, where stochastic policy (actor) and the value function (critic) are modeled by two independent neural network function approximators, i.e., a set of fully connected feed-forward networks of dimensions $256\times 128\times 64$ and the hyperbolic tangent nonlinear activation function for each hidden layer. For the nonlinear quantum optomechanical configuration, in the target-utilization phase, the recurrent PPO agent outperforms the PPO agent, where both independent critic and actor networks are MLPs followed by one independent layer of LSTM with $256\times 128\times 64$ fully connected networks and $256$ hidden states. More details are described in Supplementary Note 6.
Data availability
The data generated in this study about training results can be found in this Zenodo repository: https://doi.org/10.5281/zenodo.12584159 [101].
Code availability
The codes used in this paper can be found in the repository:
https://github.com/liliyequantum/Entanglement-engineering-by-RL [102].
Supplementary information
Supplementary information for this study is available on GitHub:
https://github.com/liliyequantum/Entanglement-engineering-by-RL [102].